You bowl 3 games and get an average score of 120. How many additional consecutive games with scores of 145 would you need to bring your average up to 130?
Answer by unlockmath(1688)

Hello,
Let's have x be the number of games:
We know 3 games average 120 so we can write the following equation:
3(120)+ 145x/3+x=130
Multiply by 3+x to get:
360+145x=390+130x
Subtract 130x and 360 to get:
15x=30
x=2
So we know it will take two more games of 145.
Make sense?
RJ
www.math-unlock.com